Project

General

Profile

Feature #1965

D1 Field Measurement

Added by Nebrensky, Henry about 4 years ago. Updated about 4 years ago.

Status:
Closed
Priority:
Normal
Category:
Beam Line
Start date:
12 April 2018
Due date:
28 May 2018
% Done:

100%

Estimated time:

Description

We would like to measure the current -> field-strength relationship as seen by the particles in traversing D1.
This can be done by removing the Q123-D1 section of the beampipe to allow access inside D1, and measuring the field as the current is ramped. This could be done during the ISIS short shut-down in late May 2018.

The purpose of this ticket is to co-ordinate this activity: those interested (or not) in taking part should add (or remove) themselves as "watchers".

The exercise will need the following to be working
  • Dipole controls (EPICS)
  • Monitoring of cooling water circuits and polarity sensors
  • Trench cooling water & Loading Bay chiller
  • PPS for permit (Muon Test)
  • Gaussmeter
  • Radiation survey (ISIS Health Physics)
  • Target phone
Personnel required:
  • 1x magnet controls operator (MLCR)
  • 2x synch team: measurement taker and monitor
  • 1x extra person in case of illness

Based on the RADOS readings last maintenance day (2┬ÁSv in 10 minutes) I've estimated that the synch team should have 4 hours to make measurements - this is tight for covering the full current range with two separate probes, so may need to consider having a second team.
(This is an under-estimate as we'd have had several days for things to cool, but we have to go by conditions at the time.)


Files

difference.png (13.6 KB) difference.png Franchini, Paolo, 26 April 2018 18:05
#1

Updated by Nebrensky, Henry about 4 years ago

The Q123-D1 section of pipe was removed during the Easter break: see photos

The wall thickness of the beam chest/box is about 1/2" and the clear aperture is about 5" (this is consistent with the nominal pole spacing of 6"), implying we need a puck about 2.5" / 6cm high.

Note the extra restrictions as the circular flange overlaps the square section:

#2

Updated by Nebrensky, Henry about 4 years ago

ISIS have confirmed they expect to have cooling water for D1 available through next shutdown. We should run later in the shutdown when the synch has cooled.
I am taking Wed. to Friday (23rd to 25th May) as the window of opportunity, intending to take the measurements themselves on Thursday 24th but with the expectation this will move should ISIS need to take priority. We'd then have Health Physics in on Wed. and have Friday to replace the beampipe.

The work will entail entry into an ISIS Controlled Area for ionising radiation, so everybody MUST have
  • a current site safety induction (Monday mornings), and
  • a current ISIS user safety test

(see RAL_ISIS_Induction - if they have expired (or you're not sure) contact to arrange)

You will also need
  • a swipe badge with R5.4 RF Hall access
  • a radiation badge, noting the System of Work for the synchrotron
  • on the day, a RADOS.

I would leave these until May, as they'll expire on 30th April if done immediately.

#3

Updated by Franchini, Paolo about 4 years ago

The Nikhef Hall probe (ID 66) shows very little discrepancies wrt the Hirst Gaussmeter (short probe).

#4

Updated by Nebrensky, Henry about 4 years ago

Hi all,
Please note that anybody needing synch access will need swipe access to R5.4 RF Hall, which needs approval from ISIS; anyone who doesn't have this should therefore fill out the user part of the permission form and send it back to me ASAP. All those I've received by end of the 15th May I will try to get signed off en masse.
If in doubt, it's easier to just fill the form in again...

#5

Updated by Nebrensky, Henry about 4 years ago

I've had the access forms for Craig and Tom signed off by ISIS.

John Govans and I have run the Trench water for half an hour - everything looked OK and no sign of major leaks. John will run up both the Trench water and the Loading-Bay Chiller circuits on Monday, so all controls / monitoring should be active from then until the end of the week.

#6

Updated by Kurup, Ajit about 4 years ago

I've had a look at the controls software and all appears to be working. The controls phone number was disconnected so I have a new number now, it's 07546432914. Give me a call if there are any issues.

#7

Updated by Franchini, Paolo about 4 years ago

Hi Henry,

have you already asked ISIS when the synch would be open and cleared for us? Just to let Craig and Tom know when they should be at RAL,

cheers,

Paolo

#8

Updated by Nebrensky, Henry about 4 years ago

Rats - I got distracted from this on Wednesday :(.

I think Health Physics were expecting to do the radiation survey on the Tuesday, so if those numbers are OK we can either go for for Wednesday while the iron's hot or Thursday as planned.

As above, my expectation is that people will be available for Wed. through Friday: we still need to be able to react to changing circumstances.

#9

Updated by Nebrensky, Henry about 4 years ago

It turned out that there'd been some misunderstanding and Health Physics were planning to do the survey on Wednesday. I asked and they've kindly done the surveys of both the synch and DSA (D2) already, with no reported concerns (though I haven't seen the numbers yet). The MICE Hall water has been conditioned and I will check the power supplies this afternoon, but it looks like the equipment will be ready to start first thing tomorrow if we are ready.

#10

Updated by Franchini, Paolo about 4 years ago

  • Status changed from Open to In Progress
  • % Done changed from 0 to 50

The HP measurements from today can be read and calibrated so I do not think we need any other measurement tomorrow. Thanks to Craig, Henry and Tom.

#11

Updated by Nebrensky, Henry about 4 years ago

  • Status changed from In Progress to Closed
  • % Done changed from 50 to 100

So the final data-taking has been completed a day ahead of schedule!

Next week we will return the power supply keys and mothball the conventional magnets.

Also available in: Atom PDF